OpenStudy (linda):

solve for x: 4x over x-3 = 2 + 12 over x-3

OpenStudy (anonymous):

(4x/x-3)=(2+12/x-3) key: *=multiply /=divide (x-3)*(4x/x-3)=(2+12/x-3)*(x-3) 4x=2+12 4x=14 (4x/4)=(14/4) x=7/2

OpenStudy (linda):

thanks!